OpenStudy (anonymous):

Marylou is blowing up a beach ball. The package says that the ball has a diameter of 12 inches. What is the volume of the beach ball when it is completely full? Use 3.14 for . A. 452.16 cubic inches B. 7,234.56 cubic inches C. 301.44 cubic inches D. 904.32 cubic inches

OpenStudy (anonymous):

@pinklion23

OpenStudy (pinklion23):

here

OpenStudy (pinklion23):

okay do you know what it is

OpenStudy (anonymous):

I think its C

OpenStudy (pinklion23):

A i believe is the answer but hold on

OpenStudy (anonymous):

\(\sf V = \dfrac{4}{3} \pi r^3\)

OpenStudy (pinklion23):

A=4πr^2 so i think it is 452.39

OpenStudy (anonymous):

plug in wht we kno: \(\sf V = \dfrac{4}{3}(3.14)(6^3)\)

OpenStudy (anonymous):

@pinklion23 ur not supposed to giv answers

OpenStudy (anonymous):

and that's wrong, it's v = 4/3pi r^3

OpenStudy (pinklion23):

im am not i was helping this person for the past 3 question we were find :)

OpenStudy (anonymous):

@ann___ can u multiply 4/3 * 3.14 * 6 * 6 * 6?

OpenStudy (anonymous):

wait im so confused i only asked @pinklion23 to help me

OpenStudy (anonymous):

some ppl cum and help no mattr what

OpenStudy (anonymous):

i'm one of them

OpenStudy (anonymous):

@ann___ so..can u multiply 4/3 * 3.14 * 6 * 6 * 6?

OpenStudy (anonymous):

that will giv u the answr

OpenStudy (pinklion23):

okay well @ann___ do you still have any other questions? like i said i think it is A

OpenStudy (anonymous):

@pinklion23 i told u ur not supposed to give answers..and that's wrong

OpenStudy (pinklion23):

okay well ask google :) https://www.google.com/search?q=how+to+find+area+od+sphere&ie=utf-8&oe=utf-8#q=how+to+find+area+of+sphere there i just put that in see what you get :)

OpenStudy (anonymous):

its volume not area, and she can't use google when she does test

OpenStudy (pinklion23):

i know why she can't but you can, i just wanted to check MY answer after you said yours that is all :)

OpenStudy (pinklion23):

so ann you still need help ? :)

OpenStudy (anonymous):

yes

OpenStudy (pinklion23):

okay so what do you think it is again :)

OpenStudy (pinklion23):

thx for the medal too lol :D

OpenStudy (pinklion23):

okay so can you tell me why you think it is C :)